OpenStudy (anonymous):

A circle has an area of 26 pie centimeters. What is the diameter? A) 6 cm B) 1,296π cm C) 12 cm D) 12π cm

OpenStudy (binary3i):

area must be pi 26 cm^2. area = pi (radius)^2 26 pi = pi R^2 where r is radius. 26=R^2 R= sqrt{26} diameter = 2R= 2sqrt{26}